OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

solve first equation for y and put it in the second equation

OpenStudy (tennistar):

8x + 2y = 16 x - 2y = 2----------ADD 9x = 18 x = 2-----------Option 2 8 + y = 8 y = 0

OpenStudy (anonymous):

is it 2 ?

OpenStudy (tennistar):

yes